OpenStudy (anonymous):

Divide. Use synthetic division, if possible. (x^4+6x^3-x^2-5x+1)/(x-2) (x^3+9x^2-5x+11)/(x^2+2)

OpenStudy (anonymous):

\[\frac{(x^4+6x^3-x^2-5x+1)}{(x-2)}\] First you would want to factor out from the first 2 terms such that x-2 would be left out, for that we can write \[6x^3=8x^3-2x^3\] \[\implies \frac{(x^4-2x^3+8x^3-x^2-5x+1)}{(x-2)}\] Take out x cube\[\frac{(x^3(x-2)+8x^3-x^2-5x+1)}{(x-2)}\] Now try to write -x^2 as something so that when u take out 8x^2 common from 8x^3 and -x^2 you would be left with x-2 just like we did with first 2 terms

OpenStudy (anonymous):

Isn't synthetic division where you only use the coefficients? i guess im having more trouble doing with the second one-dividing the x^2+2

OpenStudy (anonymous):

I don't know what's "synthetic" division, but u can also simply just divide it, but I prefer to take out factors like that, I find it simpler

OpenStudy (anonymous):

Have u done the first one?

OpenStudy (anonymous):

you know it would be so much easier to do it on calcilator

OpenStudy (anonymous):

i think synthetic division is the one where you have the divisor boxed in the top left

OpenStudy (anonymous):

i forget if u box a positive 2 or a negative 2

OpenStudy (anonymous):

Like this? 

OpenStudy (anonymous):